# SKD-TSTSAN: Three-Stream Temporal-Shift Attention Network Based on Self-Knowledge Distillation for Micro-Expression Recognition

### v1

@@ 4.5.1 Ablation Studies
We introduce a channel attention mechanism by incorporating ECA modules into the L-stream of the network, which improves UF1 and UAR by 1.13% and 0.12% on the CASME II dataset, and 2.48% and 3.89% on the MMEW dataset. Furthermore, to better understand how ECA modules influence the final recognition performance, the gradient-weighted class activation mapping (Grad-CAM) [50] is utilized for visualization analysis. Fig. 7(a), (b), and (c) represent the samples of three different emotion categories in the CASME II dataset: “Negative (Disgust)”, “Positive (Happiness)”, and “Surprise”. The apex frames are shown in the first row, and the class activation maps (CAMs) of the SKD-TSTSAN with and without the ECA module are shown in the second and third rows. These maps are from the fourth convolutional layer of the S-stream, and the red regions highlight the areas that are crucial for the MER results. According to the Facial Action Coding
@@
System (FACS), the “Disgust” expression usually involves AU9 (“Nose Wrinkler”), AU10 (“Upper Lip Raiser”), and AU4 (“Brow Lowerer”) with AU7 (“Lid Tightener”). The “Happiness” expression often involves AU6 (“Cheek Raiser”) and AU12 (“Lip Corner Puller”). The “Surprise” expression is usually found in AU5 (“Upper Lid Raiser”), AU26 (“Jaw Drop”), AU1 (“Inner Brow Raiser”), and AU2 (“Outer Brow Raiser”). As seen in the CAMs in Fig. 7(a), (b), and (c), when ECA modules are used, the facial regions with strong activation are consistent with the AUs defined by FACS. For instance, in Fig. 7(a), the eyebrow region is highlighted; in Fig. 7(b), the cheek and lip corner regions are emphasized; in Fig. 7(c), the eyebrow and chin regions are highlighted. However, when the ECA module is not used, the model tends to focus on regions less relevant to the expressions. For example, the model without ECA modules focuses on the cheek area in Fig. 7(c), which is not relevant to the “Surprise” expression. Therefore, by incorporating ECA modules into the L-stream, we not only suppress redundant information in the L-stream input but also enable the S-stream to focus more on the key
@@
regions of ME muscle movements.
Next, the introduction of TSM in the T-stream enables the exchange of temporal information between adjacent frames and also improves the performance on CASME II and MMEW datasets. When both ECA modules and TSMs are used, the performance is better than using only one of them, demonstrating the complementarity of the introduced modules. Furthermore, when the self-knowledge distillation (SKD) is introduced into MER, UF1 and UAR on CASME II are improved by 1.54% and 0.37%, and UF1 and UAR on MMEW are increased by 3.33% and 3.46%, respectively. For an intuitive analysis of the effect of SKD, as shown in Fig. 8, t-distributed stochastic neighbor embedding (t-SNE) is employed to visualize the learned ME features of subjects 1 and 12 in the CASME II dataset. Fig. 8(a) and (b) illustrate the embedding ME features of the network without and with SKD, respectively. It can be easily observed that with the addition of SKD, the clustering within the same category is closer, and the decision boundaries between different categories are clearer, thus the network has better performance. In addition, the deepest section of the network serves as the teacher model in SKD without the need to design and train an additional teacher model, which reduces the training time.
@@ 4.5.3 More Analysis on the Choice of Amplification Factors
To evaluate the effect of different amplification factors (AFs) on the performance, we provide the UAR under different AF configurations on the SAMM and MMEW datasets in Fig. 9(b). For the 3-class MER task on the SAMM dataset, our SKD-TSTSAN consistently outperforms the others MER methods ($\leq$81.24UAR) regardless of the AF value, indicating the robustness and effectiveness of our approach. When the AF is too small, the facial muscle movements are not obvious; while when the AF is too large, the muscle movements are excessively deformed, both of which affect the MER performance. When the AF is set to 2, the results are optimal on both the SAMM and MMEW datasets, so this value is used as our final configuration.
@@
4.5.4 More Analysis on the Temporal Domain Distribution
As illustrated in Fig. 10(a), in our comparative experiment configuration, our T-stream inputs are optical flow maps from two different temporal domains, calculated from the onset frame, apex frame, and offset frame. To explore the effect of different temporal distributions of T-stream inputs, as shown in Fig. 10(b) and (c), we first interpolate each ME sequence to 6 and 11 frames using the temporal interpolation model (TIM) [5]. The optical flow maps are then calculated between adjacent frames, representing the motion information from 5 and 10 different temporal domains, respectively. Subsequently, the optical flow maps from different temporal domains are input into the T-stream, where temporal modeling is performed using TSM. Table 6 provides the performance of our SKD-TSTSAN on the CASME II dataset under different temporal domain distributions. It’s clear that the optimal results are obtained when we use the optical flow between the three key-frames as T-stream input, with an improvement of 2.96% for UF1 and 2.80%
for UAR compared to inputting 10 optical flow maps. The main reason is that the onset frame and offset frame represent the start and end of ME, while the apex frame represents the peak intensity of the expression. The optical flow maps between the three key-frames effectively reflect the motion information. After interpolation using TIM, the motion amplitude between adjacent frames is smaller, resulting in less distinct optical flow features and worse results.
@@ 4.5.5 More Analysis on the Performance of Different Classifiers
To verify that deeper classifiers in SKD extract more discriminative features, we compare the performance of each classifier on multiple datasets. As shown in Table 7, the deepest classifier 3/3 achieves the highest UF1 and UAR scores on different datasets. The deeper classifier 2/3 performs better than the shallow classifier 1/3, since the features extracted from the deeper
@@
section are more discriminative. Although classifier 2/3 does not perform as well as the deepest classifier 3/3, its performance can also outperform other MER methods on different datasets, indicating the effectiveness of our SKD-TSTSAN. To visualize the discriminative features extracted by different classifiers, t-SNE is utilized to visualize the embedding ME features of subjects spNO.142 and spNO.176 in the CAS(ME)3 dataset, as shown in Fig. 11. It is obvious that the deeper classifiers have more concentrated clustering of the same category and clearer decision boundaries between different categories, indicating that more discriminative features are extracted.
@@ 5 Conclusion
In this paper, we classify MEs into a set of emotion categories using the onset frames, apex frames, and offset frames labeled in the dataset. However, to achieve an end-to-end ME analysis system, we still need to design an ME spotting method to identify key-frames. Therefore, in future work, we will explore ME spotting methods in long videos. Additionally, many current research efforts treat the ME spotting and recognition tasks separately [11]. However, these two tasks are strongly related and both require extracting subtle facial muscle motion features. Therefore, we will attempt to integrate ME spotting with our designed ME recognition network to implement ME analysis in a unified network.

### v2

@@ 4.5.1 Ablation Studies
We introduce a channel attention mechanism by incorporating ECA modules into the L-stream of the network, which improves UF1 and UAR by 1.13% and 0.12% on the CASME II dataset, and 2.48% and 3.89% on the MMEW dataset. Furthermore, to better understand how ECA modules influence the final recognition performance, the gradient-weighted class activation mapping (Grad-CAM) (Selvaraju et al.,, 2017) is utilized for visualization analysis. Fig. 7(a), (b), and (c) represent the samples of three different emotion categories in the CASME II dataset: “Negative (Disgust)”, “Positive (Happiness)”, and “Surprise”. The apex frames are shown in the first row, and the class activation maps (CAMs) of the SKD-TSTSAN with and without the ECA module are shown in the second and third rows. These maps are from the fourth convolutional layer of the S-stream, and the red regions highlight the areas that are crucial for the MER results. According to the Facial Action Coding System (FACS), the “Disgust” expression usually involves AU9 (“Nose Wrinkler”), AU10 (“Upper Lip Raiser”), and AU4 (“Brow Lowerer”) with AU7 (“Lid Tightener”). The “Happiness” expression often involves AU6 (“Cheek Raiser”) and AU12 (“Lip Corner Puller”). The “Surprise” expression is usually found in AU5 (“Upper Lid Raiser”), AU26 (“Jaw Drop”), AU1 (“Inner Brow Raiser”), and AU2 (“Outer Brow Raiser”). As seen in the CAMs in Fig. 7(a), (b), and (c), when ECA modules are used, the facial regions with strong activation are consistent with the AUs defined by FACS. For instance, in Fig. 7(a), the eyebrow region is highlighted; in Fig. 7(b), the cheek and lip corner regions are emphasized; in Fig. 7(c), the eyebrow and chin regions are highlighted. However, when the ECA module is not used, the model tends to focus on regions less relevant to the expressions. For example, the model without ECA modules focuses on the cheek area in Fig. 7(c), which is not relevant to the “Surprise” expression. Therefore, by incorporating ECA modules into the L-stream, we not only suppress redundant information in the L-stream input but also enable the S-stream to focus
@@
more on the key regions of ME muscle movements.
Next, the introduction of TSM in the T-stream enables the exchange of temporal information between adjacent frames and also improves the performance on CASME II and MMEW datasets. When both ECA modules and TSMs are used, the performance is better than using only one of them, demonstrating the complementarity of the introduced modules. Furthermore, when the self-knowledge distillation (SKD) is introduced into MER, UF1 and UAR on CASME II are improved by 1.54% and 0.37%, and UF1 and UAR on MMEW are increased by 3.33% and 3.46%, respectively. For an intuitive analysis of the effect of SKD, as shown in Fig. 8, t-distributed stochastic neighbor embedding (t-SNE) is employed to visualize the learned ME features of subjects 1 and 12 in the CASME II dataset. Fig. 8(a) and (b) illustrate the embedding ME features of the network without and with SKD, respectively. It can be easily observed that with the addition of SKD, the clustering within the same category is closer, and the decision boundaries between different categories are clearer, thus the network has better
@@
performance. In addition, the deepest section of the network serves as the teacher model in SKD without the need to design and train an additional teacher model, which reduces the training time.
@@ 4.5.3 More Analysis on the Choice of Amplification Factors
To evaluate the effect of different amplification factors (AFs) on the performance, we provide the UAR under different AF configurations on the SAMM and MMEW datasets in Fig. 9(b). For the 3-class MER task on the SAMM dataset, our SKD-TSTSAN consistently outperforms the others MER methods ($\leq$81.24UAR) regardless of the AF value, indicating the
robustness and effectiveness of our approach. When the AF is too small, the facial muscle movements are not obvious; while when the AF is too large, the muscle movements are excessively deformed, both of which affect the MER performance. When the AF is set to 2, the results are optimal on both the SAMM and MMEW datasets, so this value is used as our final configuration.
4.5.4 More Analysis on the Temporal Domain Distribution
As illustrated in Fig. 10(a), in our comparative experiment configuration, our T-stream inputs are optical flow maps from two different temporal domains, calculated from the onset frame, apex frame, and offset frame. To explore the effect of different temporal distributions of T-stream inputs, as shown in Fig. 10(b) and (c), we first interpolate each ME sequence to 6 and 11 frames using the temporal interpolation model (TIM) (Pfister et al.,, 2011). The optical flow maps are then calculated between adjacent frames, representing the motion information from 5 and 10 different temporal domains, respectively. Subsequently, the optical flow maps from different temporal domains are input into the T-stream, where temporal modeling is performed using TSM. Table 6 provides the performance of our SKD-TSTSAN on the CASME II dataset under different temporal domain distributions. It’s clear that the optimal results are obtained when we use the optical flow between the three key-frames as T-stream input, with an improvement of 2.96% for UF1 and 2.80% for UAR compared to inputting 10 optical flow maps. The
@@ 4.5.4 More Analysis on the Temporal Domain Distribution
main reason is that the onset frame and offset frame represent the start and end of ME, while the apex frame represents the peak intensity of the expression. The optical flow maps between the three key-frames effectively reflect the motion information. After interpolation using TIM, the motion amplitude between adjacent frames is smaller, resulting in less distinct optical flow features and worse results.
@@ 4.5.5 More Analysis on the Performance of Different Classifiers
To verify that deeper classifiers in SKD extract more discriminative features, we compare the performance of each classifier on multiple datasets. As shown in Table 7, the deepest classifier 3/3 achieves the highest UF1 and UAR scores on different datasets. The deeper classifier 2/3 performs better than the shallow classifier 1/3, since the features extracted from the deeper section are more discriminative. Although classifier 2/3 does not perform as well as the deepest classifier 3/3, its performance can also outperform
@@
other MER methods on different datasets, indicating the effectiveness of our SKD-TSTSAN. To visualize the discriminative features extracted by different classifiers, t-SNE is utilized to visualize the embedding ME features of subjects spNO.142 and spNO.176 in the CAS(ME)3 dataset, as shown in Fig. 11. It is obvious that the deeper classifiers have more concentrated clustering of the same category and clearer decision boundaries between different categories, indicating that more discriminative features are extracted.
@@ 5 Conclusion
In this paper, we classify MEs into a set of emotion categories using the onset frames, apex frames, and offset frames labeled in the dataset. However, to achieve an end-to-end ME analysis system, we still need to design an ME spotting method to identify key-frames. Therefore, in future work, we
will explore ME spotting methods in long videos. Additionally, many current research efforts treat the ME spotting and recognition tasks separately (Song et al.,, 2019). However, these two tasks are strongly related and both require extracting subtle facial muscle motion features. Therefore, we will attempt to integrate ME spotting with our designed ME recognition network to implement ME analysis in a unified network.